A BODY HAS A MASS OF 15KG AND IT IS BEING PUSHED HORIZONTALLY WITH A FORCE OF 100N IF IT HAS A COEFFICIENT OF KINETIC FRICTION 0.3 THEN WHAT IS THE ACCELERATION OF THE BODY)


1
Expert's answer
2022-01-23T16:14:11-0500

Given:

"m=15\rm\: kg"

"F=100\rm\: N"

"\mu_k=0.3"

"a-?"


The Newton's second law says

"ma=F_{\rm net}"

"F_{\rm net}=F-F_f=F-\mu_k N=F-\mu_kmg"

Hence, the acceleration of the system

"a=\frac{F_{\rm net}}{m}=\frac{F}{m}-\mu_kg"

"a=\rm\frac{100\:N}{15\:kg}-0.3*9.8\: m/s^2=3.7\: m/s^2"